platypus foraging for prey can detect an electric field as small as 0.002 N/C To give an idea of sensitivity of the platypus’s electric sense, how far from a +20nC point charge does the field have this magnitude?
Answer #1

Magnitude of electric field due to point charge q (C) at distance r (m) from it ,

E = kq/r^2 N/C

Where k = 9×10^9 N m^2/C^2

Hence for E = 0.002 N/C and q = 20 nC = 20×10^-9 C, we get

0.002 = 9×10^9 * 20×10^-9 / r^2

r = 300 m